Let V be a vector space over a filed F and x,y,z is an element of V then show that the set of all liner combinations of x,y and z,W=(ax+by+cz:a,b,c are an element of F) is a subspace of V over F. This subspace is called the span of (x,y,z)


1
Expert's answer
2021-12-09T17:29:54-0500

Let W={ax+by+cz:a,b,c "\in" "\mathbb{F}" }

We want to show that W is a subspace for x,y,z"\in" V(where V is a vector space over "\mathbb{F}" ) i.e "\alpha"u+"\beta"v"\in" W, u,v"\in" W and "\alpha" ,"\beta" "\in" "\mathbb{F}" .

Since u"\in" W => u="a_1"x + "b_1"y + "c_1"z and v"\in" W =>"a_2"x + "b_2"y + "c_2"z for"a_1,b_1,c_1,a_2,b_2,c_2 \in \mathbb{F}"

"\alpha u +" "\beta v" ="\alpha (a_1x + b_1y + c_1z)" + "\beta(a_2x +b_2y +c_2z)"

="\alpha a_1x +\alpha b_1y + \alpha c_1z +\beta a_2x +\beta b_2y+ \beta c_2z"

="(\alpha a_1 +\beta a_2)x +(\alpha b_1 +\beta b_2)y +(\alpha c_1 + \beta c_2)z"

Since ("\alpha a_1+\beta a_2),(\alpha b_1+\beta b_2),(\alpha c_1+\beta c_2) \in \mathbb{F}"

=>"px+qu+rz \in W ,p,q,r \in \mathbb{F}"

Which implies that "\alpha u + \beta v =px+qu+rz \in W ,p,q,r \in \mathbb{F}"

Hence, W is a subspace of V.